A multiple choice quiz has 400 questions, each question with five possible answers of which one is correct answer. What is the probability that a sheer guess yields exactly 80 correct answers using normal approximation.


1
Expert's answer
2021-05-31T12:29:29-0400

We have given that,


"n = 400"


"p = \dfrac{1}{5} = 0.2"


"q = 0.8"


We can say that the sampling distribution of X can be approximately normal with:


Mean "= np = 400 \times 0.2 = 80"


Standard deviation "= \sqrt{npq} = \sqrt{400\times 0.2 \times 0.8} = 8"


Now, the probability that a guesswork yields exactly "80" correct answers is :


"P(X = 80)"


"= P(79.5\le X\le80.5)"


"= 0.049"